A study found that 1 out of 200 adult males have green eyes. If a random sample of 400

adult males is​ obtained, is the sampling of the sample proportion of males that have green

eyes approximately​ normal?
Answer #1

Solution :

p = 200 / 400 = 0.5

np(1-p) 10

400 * 0.5 * 0.5 = 100

Satisfied the condition .

The sampling of the sample proportion of males that have green eyes are approximately​ normal